Having trouble on this. Any help is appreciated.
Square root of y+9 = 1

Answer by stanbon(75887)

Square root of (y+9) = 1
---------
sqrt(y+9) = 1
Square both sides to get:
y+9 =1
y = -8
===========
Cheers,
Stan H.